OpenStudy (anonymous):

2sin^2 (x)=sinx solve the equation for exact solutions over the interval [0,2pi)

OpenStudy (mathmale):

Re-writing your equation like this may make it easier for y ou and for others to decipher (understand):\[2\sin^2 (x)=sinx \rightarrow 2(\sin x)^2=\sin x\] First, think about how you would solve the following: 2c^2=c. Hint: subtract c from both sides, and then factor the resulting equation.

OpenStudy (anonymous):

Well, I was wondering, I know my answer is 0 and pi but, do you think it could be any other factors?